I feel I should know this already sorry. But why multiply by 6? (1/6)x - (2/3)x;
if you put into like terms just multiply 2/3 by 2 for ((4/6)x

Your 2nd eqn has no x term.
If you meant:
Y = (1/6)x + 3 and Y = (2/3)x + 0
----
Since they both = y,
(1/6)x + 3 = (2/3)x
Multiply by 6
x + 18 = 4x
3x = 18
x = 6
y = 4
--> (6,4)

Answer by jim_thompson5910(35256)

Fractions are generally a pain. So if they are found in equations, it's usually a good idea to clear them out.

Multiplying both sides by 6 will clear them out since 6 is the LCD (lowest common denominator)

Notice how 6*(1/6) = 6/6 = 1. So that fraction goes away
Also, notice how 6*(2/3) = 12/3 = 4. That fraction goes away as well.

-----------------------------

So (1/6)x + 3 = (2/3)x turns into x + 18 = 4x. The two equations are equivalent

Hopefully you agree that x + 18 = 4x is much simpler compared to (1/6)x + 3 = (2/3)x